OpenStudy (anonymous):

12. The owner of a store displays a large jar of nickels and dimes and offers the value of the coins to the person who guesses how many dimes there are. If there are 1,130 coins, and they are worth $100, how many dimes are there?

OpenStudy (anonymous):

5N + 10D = 100 N + D = 1130

OpenStudy (anonymous):

Can you solve from there?

OpenStudy (anonymous):

no can you tell me more details please

OpenStudy (anonymous):

Ok I see the problem sorry it should be: 0.05N + 0.10D = 100 N + D = 1130

OpenStudy (anonymous):

So, multiply eq.1 by -20 and add to eq.2 0.05N + 0.10D = 100 -2D = -870 D = 435 coins. which makes N = 695 coins.

OpenStudy (anonymous):

i got it thank you